OpenStudy (anonymous):

Graph the function f(x)=x^3−2x and it's secant line through the points (-2,-4) and (2,4). Use the graph to estimate the x-coordinate of the points where the tangent line is parallel to the secant line. Find the exact value of the numbers c that satisfy the conclusion of the mean value theorem for the interval [-2,2].

OpenStudy (anonymous):

I need help finding 'c'

OpenStudy (anonymous):

I know that the Mean Value Theorem says... \[\frac{ f(b)-f(a) }{ b-a }=f'(c)\]

OpenStudy (anonymous):

just apply that...

OpenStudy (anonymous):

But that gives me f'(c)

OpenStudy (anonymous):

lemme check

OpenStudy (anonymous):

here a=-2 b=2

OpenStudy (anonymous):

Right, plugging all that in do you get f'(c) = 2 ?

OpenStudy (anonymous):

i get 8

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

f(b)=2^3-2(2)..rite?

OpenStudy (anonymous):

f(a)=(-2)^3-2(-2)

OpenStudy (anonymous):

Doesn't it already give us the coordinates for the points? (2,4) and (-2,-4)

OpenStudy (anonymous):

Is it wrong to say \[\frac{ 4+4 }{ 2+2}=\frac{ 8 }{ 4 } = 2\] ? Or am I totally wrong right now? Lol

OpenStudy (anonymous):

what did u do just now?

myininaya (myininaya):

\[(x^3-2x)'|_{x=c}=\frac{f(2)-f(-2)}{2-(-2)} \\ (x^3-2x)'|_{x=c}=\frac{4-(-4)}{2-(-2)} \\ (x^3-2x)'|_{x=c}=\frac{4+4}{2+2}\] that is right for the right hand side

OpenStudy (anonymous):

Well I plugged in f(a) = -4, f(b) = 4, and a=-2, b=2 in the mean value theorem formula

OpenStudy (anonymous):

what happens to the left hand side?

OpenStudy (anonymous):

do i just derive it?

myininaya (myininaya):

you still have to differentiate the left hand side and plug in c afterwards as the equation above suggests

OpenStudy (anonymous):

3x^2 - 2 right?

myininaya (myininaya):

yep and plug in c (though that part doesn't matter to much; it is just what they call it in the formula above) yep now solve 3x^2-2=2

myininaya (myininaya):

or you named it c 3c^2-2=2

OpenStudy (anonymous):

Ohh so then its +/- 2/sqrt3 right?

myininaya (myininaya):

are both of those in the given interval?

OpenStudy (anonymous):

hmmm... yes?

myininaya (myininaya):

ok then cool stuff

OpenStudy (anonymous):

Thats it? :O

myininaya (myininaya):

that is it

myininaya (myininaya):

though there are two problems

myininaya (myininaya):

did you do the first problem?

OpenStudy (anonymous):

Yay!!! Thats correct! :D Thank you so much!

myininaya (myininaya):

we did the second problem we found the exact solutions

OpenStudy (anonymous):

Yeah Im always better in graphing haha

myininaya (myininaya):

alright

OpenStudy (anonymous):

Thanks again for the help! :)

myininaya (myininaya):

np

OpenStudy (anonymous):

tracy?got the answer for the second part?

OpenStudy (anonymous):

Yeah! I got it, its + or - 2/sqrt3

OpenStudy (anonymous):

Thank you too @chethus!